A car with a gun mounted on it is kept on horizontal frictionless surface. Total mass of car, gun and shell is 50kg. Mass of each shell is 1kg. If shell is fired horizontally with relative velocity 100 m/s with respect to gun. What is the recoil speed of car after second shot?

Grade:12

4 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

assume the velocity of car after recoil = -v (in opposite direction)
and the velocity of sheel wrt to ground = u
so , u – (-v) = u +v = 100
Apply conservation of momentum in this case.
49(-v) + 1(u) = 0, so , u = 49v
so, using these two equations , v = 2 m/s

thanks.
Vikrant Prasad
9 years ago
thanks sir, but the recoil velocity after second shot is asked .. but I calculated it by following the above method … Thank you :-)
Vikrant Prasad
9 years ago
its 4.04 m/s
Ranjeet yadav
6 years ago
a gun is mounted on a railroad car.the mass of the car the gun,the shell.if the muzzle velocity of the shells is 200m/s,what is the recoil speed of the car after the second shot ? neglect friction